OpenStudy (anonymous):

Solve 3x^2 - 4x - 2 = 0

OpenStudy (anonymous):

x (3 x-4) = 2 3 x^2 = 4 x+2 9/10 (x-2/3)^2 = 1 so,, x = 1/3 (2-sqrt(10)) or x = 1/3 (2+sqrt(10))

OpenStudy (anonymous):

Would it be [2 +- sqrt(10)]/3?

OpenStudy (anonymous):

2/3-sqrt(10)/3

OpenStudy (anonymous):

-0.38742